# Dynamical friction and measurements of the splashback radius in galaxy clusters

### v1

@@ 3 Results
Work by van den Bosch et al. (1999) performed N-body simulations of an isothermal halo with Milky Way-like mass $\sim 10^{12}~{}\mathrm{M_{\odot}}$, and studied eccentricities of orbiting objects. They found DF timescales to be Gyr-scale, and that DF is negligible for satellites with mass under $10^{9}~{}\mathrm{M_{\odot}}$, producing a mass fraction requirement of $M_{\mathrm{subhalo}}/M_{\mathrm{halo}}~{}>~{}0.1\%$. They argue that this 0.1% figure scales for galaxy clusters, while assuming that cluster timescales are equivalent to the Milky Way simulation timescale. Therefore they expect dynamical friction to matter only for subhaloes of total mass $>10^{12}~{}\mathrm{M_{\odot}}$, within a hypothetical cluster of mass $10^{15}~{}\mathrm{M_{\odot}}$. A subhalo mass of $10^{12}~{}\mathrm{M_{\odot}}$ corresponds to a stellar mass of $\sim 2\times 10^{10}~{}\mathrm{M_{\odot}}$ (see Table 2). Using a cut of $M_{*}>10^{10}~{}\mathrm{M_{\odot}}$ for our most massive cluster bin, we do not see a significant reduction in $R_{\mathrm{sp}}$, suggesting that a mass fraction above 0.1% may be required.
More recent work by van den Bosch et al. (2016), however, argues that a mass ratio of 10% is necessary for dynamical friction to notably alter an orbit, among haloes of mass $\gtrsim 10^{13}~{}\mathrm{M_{\odot}}$. They examine three different cosmological simulations (from the Bolshoi and Chinchilla suites) with varying cosmological parameters and mass resolutions. They find this mass fraction requirement across all three simulations. The 10% figure is consistent for our findings in the high cluster-mass regime. For example, among our most massive clusters, ($\sim 10^{15}~{}\mathrm{M_{\odot}}$) we would expect DF to be impactful only for stellar mass cuts exceeding $7\times 10^{11}~{}\mathrm{M_{\odot}}$. We thus would not expect to find significant differences for any of the stellar mass cuts shown in Figure 5, and we do not. On the other end, for a cluster of mass $10^{13}~{}\mathrm{M_{\odot}}$ in our sample, we would require a minmum stellar mass of $\sim 2\times 10^{10}~{}\mathrm{M_{\odot}}$ for DF to have an impact. We do see a significant ($\geq 1\sigma$) reduction in the splashback radius for a cut of $M_{*}>10^{10}~{}\mathrm{M_{\odot}}$ after 5 Gyr, suggesting that the requirement of van den Bosch et al. (2016) may be too stringent. However, direct comparison is challenging because they did not divide their sample by halo mass. They, additionally, account for processes such as tidal stripping that contribute to subhalo mass loss, which are excluded from our study. Hence, it is expected that a higher starting mass fraction would be required if mass loss processes are included, because mass loss should reduce the strength of dynamical friction.
Yet other studies have found that a mass fraction of 1% is the requirement for dynamical friction to have an impact on subhaloes (Adhikari et al., 2016; More et al., 2016). Adhikari et al. (2016) examined clusters of virial mass $\geq 10^{14}$, and subhaloes with radial plunging orbits. They found the splashback radius was reduced by up to $\sim 20\%$ for low-mass clusters ($3-5\times 10^{13}~{}\mathrm{M_{\odot}}$) between subhaloes with masses less than- versus exceeding $4\times 10^{11}~{}\mathrm{M_{\odot}}$, which corresponds to a stellar mass cut of $\sim 3\times 10^{9}~{}\mathrm{M_{\odot}}$. In the matching cluster mass bin ($10^{13.5}\leq M_{\mathrm{200,mean}}/\mathrm{M_{\odot}}<10^{14}$), we find only a 7% reduction in the splashback radius, even using a stellar mass cut of $10^{10}~{}\mathrm{M_{\odot}}$. We do not find as significant a reduction as Adhikari et al. (2016) for that case. Plunging radial orbits (as used in Adhikari et al. (2016)) should introduce the most extreme case of DF, so it is unsurprising that their subhaloes are more affected than those of this simulation. They also find that DF has a reduced effect in lower-concentration clusters, bringing their results closer to ours. Future work could examine the role of cluster concentration with the IllustrisTNG sample, using our more physically motivated subhalo orbits.
The wide variation in literature mass fractions reinforces the need for numerical simulations to ascertain the complex dependence of DF on subhalo mass, size, position, velocity, as well as halo density. We generally find that requiring $M_{\mathrm{subhalo}}/M_{\mathrm{halo}}~{}>~{}1\%$ captures our results, though again, we find that $M_{\mathrm{*}}/M_{\mathrm{halo}}$ is a more useful metric. For clusters of mass greater than $10^{14}\mathrm{M_{\odot}}$, a reduced splashback radius should be traced only by galaxies with stellar masses exceeding $10^{10}$ M, consistent with Figure 5.
@@
Though these variations in $R_{\mathrm{sp}}$ are interesting, we are primarily interested in the discrepancies between the simulations run with and without dynamical friction. To this end, Figure 6 shows the ratio of the splashback radii with and without dynamical friction, but as a function of minimum subhalo stellar mass over halo mass. By looking at the overlapping points, we can see that despite the residual trends seen in Figs. 4 and 5, there is very good agreement when we examine the fraction $R_{\mathrm{sp,DF}}/R_{\mathrm{sp,NoDF}}$. Therefore we conclude that any contamination of our sample does not significantly affect our analysis of dynamical friction.
It is only after this mass fraction ($M_{*}/M_{\mathrm{200,mean}}$) exceeds $10^{-4}$ that we see notable decreases in the splashback radius ($\sim 10\%$) due to DF, which is consistent with our estimates and the figure produced by Adhikari et al. (2016). We also note that, as expected, the mass fraction is what drives the trend here, not only the subhalo mass cut or cluster mass. However, O’Neil et al. (2021) showed that in the TNG simulations it was the absolute subhalo mass that matters for the decrease in splashback radius, again indicating that dynamical friction cannot be the only reason for those offsets. This difference is particularly of note since we use the same sample as O’Neil et al. (2021).

### v2

@@ 3 Results
Though these variations in $R_{\mathrm{sp}}$ are interesting, in this study, we are primarily interested in the discrepancies between the simulations run with and without dynamical friction. To this end, Figure 6 shows the ratio of the splashback radii with and without dynamical friction, but as a function of minimum subhalo stellar mass over halo mass. By looking at the overlapping points, we can see that despite the residual trends seen in Figs. 4 and 5, there is very good agreement when we examine the fraction $R_{\mathrm{sp,DF}}/R_{\mathrm{sp,NoDF}}$. Therefore we conclude that any contamination of our sample does not significantly affect our analysis of dynamical friction.
It is only after this mass fraction ($M_{*}/M_{\mathrm{200,mean}}$) exceeds $10^{-4}$ that we see notable decreases in the splashback radius ($\lesssim 10\%$) due to DF, which is consistent with our estimates and the figure produced by Adhikari et al. (2016). We also note that, as expected, the mass fraction is what drives the trend here, not only the subhalo mass cut or cluster mass. However, O’Neil et al. (2021) showed that in the TNG simulations it was the absolute subhalo mass that matters for the decrease in splashback radius, again indicating that dynamical friction cannot be the only reason for those offsets. This difference is particularly of note since we use the same sample as O’Neil et al. (2021).
@@
3.1 Residence Time
The splashback radius is traced out by subhaloes that have spent sufficient time in the cluster to reach their first apocenter. In position-radial velocity phase-space, then, $R_{\mathrm{sp}}$ should appear as an overdensity of subhaloes with low radial velocities. Moreover, galaxies that are just infalling into the cluster should not trace out the splashback radius. In this section, we investigate how the cluster phase-space evolves with residence time in our idealized simulations.
The phase-space distributions for high-stellar mass galaxies ($M_{*}>10^{9}~\mathrm{M_{\odot}}$) are shown in Figure 7 after 10 Gyr of evolution, for our lowest-mass clusters (between $10^{13}$ and $10^{13.5}\mathrm{M_{\odot}}$). We examine the phase space after 10 Gyr, rather than 5 Gyr like in the rest of the analysis, to see longer-duration changes in the cluster environment. We separate the subhaloes by residence time, which is defined as the time since the subhalo first crossed inside of $R_{\mathrm{200,mean}}$. The top three panels show subhaloes with residence times of 1 to 3 Gyr, 4 to 6 Gyr, and 7 to 10 Gyr. The bottom panel shows all galaxies meeting the mass cut, including galaxies that never reach within $R_{200,\mathrm{mean}}$ or galaxies that began the simulation inside $R_{200,\mathrm{mean}}$. We filter out a small number of galaxies with $M_{\mathrm{subhalo}}>0.05M_{200,\mathrm{mean}}$, as they should, in a physically realistic environment, influence the movement of the cluster. On the plot they would appear as a horizontal band throughout a range of radial positions with very low radial velocities (approximately zero).
If instead a minimum stellar mass cut of $10^{7}~\mathrm{M}_{\odot}$ is applied, the phase-space profiles resemble the expected distribution from IllustrisTNG outputs (see e.g. Dacunha et al., 2022; O’Neil et al., 2024), suggesting our model reasonably recovers the phase-space measured by other simulated clusters (e.g. the large number of infalling galaxies with negative radial velocity, the dispersal of galaxies at long residence times), even after 10 Gyr of idealized simulation. For galaxies with residence times between 1 and 3 Gyr, we see the infalling and outgoing galaxies combined, as well as a turnaround in phase space around $\sim 5$ Gyr. Thus, we conclude that our simplified clusters (e.g. treating subhaloes as point masses) maintain the expected phase space from N-body simulations.
Figure 6: The ratio of the splashback radius with and without dynamical friction, after 5 Gyr, for the ratio of minimum stellar mass to halo mass (measured in the middle of the bin). The horizontal grey dotted line marks a ratio of 1 (i.e. no difference). The vertical black dashed line shows where $M_{\mathrm{subhalo}}/M_{\mathrm{200,mean}}\approx 0.01$. Finally, the solid grey line and shaded band show average ratio and error for the regions with multiple data points.
The splashback feature is thought to be found among galaxies with residence times of order one crossing time ($\sim$ 5 Gyr, from O’Neil et al., 2024), with radial velocities around 0. We begin to see these tracer galaxies in the middle residence time bin, and the number grows with residence times of over 7 Gyr (the top panel). While one might expect the splashback feature to be traced by galaxies with residence times of closer to 5 Gyr, the feature is in fact also visible among galaxies that have spent longer in the cluster (residence time between 7 and 9 Gyr). We attribute this phenomenon to weaker gravitational forces in the least-massive clusters producing less plunging orbits.
Figure 7: 2D histograms of radial velocity and radius plots after 10 Gyr of evolution, with (right column) and without (left) dynamical friction, for the lowest mass haloes ($10^{13}$ - $10^{13.5}$ M, totaling 770 haloes) and subhaloes with minimum stellar mass $M_{*}~=~10^{9}$ M. Subhaloes with $M_{\rm subhalo}~>~0.05M_{200,\mathrm{mean}}$ are also excluded. The top three panels are separated by residence time in the cluster, which is calculated from the first time at which the galaxy reaches a position within $R_{200,\mathrm{mean}}$. The bottom row shows the histogram for all subhaloes with $M_{*}$ > 109 M and $M_{\mathrm{subhalo}}~<~0.05~M_{200,\mathrm{mean}}$, totaling over 47,000 galaxies. The black dashed line shows the median splashback radius with and without dynamical friction (as measured in Figure 5.)
The majority of galaxies in this halo sample never reach inside of $R_{200,\mathrm{mean}}$ (approximately 88% whether with or without DF.) Of the galaxies that do, the influence of dynamical friction is most apparent among the galaxies with longer residence times. There are few differences between the phase-space distributions for galaxies with $t_{\mathrm{res}}<4$ Gyr. By the longest residence time bin, though, the splashback radius (marked with a vertical dashed line) is slightly reduced. The distribution of radii at $v_{r}\approx 0$ is wider with dynamical friction enabled as well, matching the wider bootstrapped $R_{\mathrm{sp}}$ distributions shown in Figure 8.
A small number of galaxies with $v_{r}<0$ and $4\leq t_{\mathrm{res}}<7$ appear only with dynamical friction enabled. By comparison to the phase-space distribution with a lower mass cut, it is apparent that (essentially) only galaxies with high masses exist in this region of phase space, and only when dynamical friction is acting. We can thus see how DF, for a small number of high mass galaxies, dramatically affects their orbital properties (e.g. as visualized in Fig. 1). However, despite these cases, it is also clear that the phase-space distributions are not, overall, dramatically altered with dynamical friction enabled.
4 Discussion
This work is motivated by a lack of studies specifically targeting the role of dynamical friction in the galaxy cluster context. As far as the authors are aware, only Adhikari et al. (2016) specifically isolates DF in their toy model simulations, finding that DF reduced the splashback radius, when measured through galaxy number densities, in a spherical system where all galaxies moved in perfectly radial orbits. Adhikari et al. (2016) concluded not only that dynamical friction resulted in a noticeable decrease in $R_{\mathrm{sp}}$ for subhaloes above 1% of the host halo mass, but also that the influence of DF is dependent on the cluster potential profile. We are motivated to create a complementary study.
4.1 Limiting assumptions
Practically speaking, isolating DF requires numerous simplifying assumptions about the cluster environment, which forms and evolves through the intersection of many processes (outlined briefly in the Introduction). The initial conditions of our simulation (i.e. the $z=0$ TNG snapshot) evolved with dynamical friction, thus impacting the orbits of the galaxies we evolve further. Although we allow one crossing time to pass before analyzing the location of $R_{\rm sp}$, we cannot fully quantify the impact of our initial conditions on our findings, whether that might produce an overestimate (e.g., due to pre-existing mass-segregation) or underestimate in the role of DF. However, as both the with- and without- DF simulations start from identical snapshots, we believe it is still useful to compare between the runs. We stress that our results, therefore, do not aim to accurately calculate values of the splashback radius, but instead to gain a preliminary upper limit on how $R_{\mathrm{sp}}$ measurements might be affected by the presence of dynamical friction. This includes our use of a static spherical gravitational potential for the clusters and the lack of baryonic physics applied to the subhaloes – each of which we discuss in the Methods section. More complexity (e.g. the inclusion of baryonic physics) could be added in future studies, as this work is only a first step towards more rigorously understanding processes that impact measurements of the splashback radius and other proxies that rely on the galaxy number density profile.
The use of a gravitational potential rather than N-body calculations is necessary in order to avoid creating dynamical friction. We also intentionally exclude accretion in order to isolate dynamical friction. The unmodeled growth of the halo potential with time may cause an underestimate of DF. However, work by Diemer et al. (2013); Wu et al. (2013) indicates that most of the cluster mass growth from $z=1$ (>5>5 Gyr ago) until the present day is due to pseudoevolution, rather than physical accretion of mass. This minimal real mass accretion should be particularly negligible and continue into the simulated future for our sample of relaxed, isolated clusters.
We treat our subhaloes as point masses with fixed mass traversing a spherical gravitational potential. Both of these choices are not strictly accurate, as clusters are elliptical and galaxies lose mass through stripping as they enter the cluster. Nevertheless, both of these choices are consistent with the methodology used in Adhikari et al. (2016). The treatment of clusters as spherically symmetric follows that of Adhikari et al. (2014), who find that the use of spherical cluster potentials does not substantially alter their measurement of the splashback radius (see also Olamaie et al., 2012, for further discussion of sphericality of clusters). Furthermore, whether observed or simulated, all downstream calculations of the differential density profile in the literature assume a spherical halo profile, consistent with our methodology. In fact, as the subhalo velocities and properties are drawn from TNG data, the subhalo phase space, and the direction of infall, is improved relative to the radial-only movement of the subhaloes in (Adhikari et al., 2016).
Our exclusion of mass stripping is necessary to reduce confounding factors in our study of dynamical friction alone. As both ram-pressure and tidal stripping are dependent on the distance from the cluster center, subhaloes may experience different mass loss rates if their orbits are decayed by DF. In addition to potentially confounding our results, incorporation of baryons into our simulations would significantly expand the scope and computational requirements of this work, without commensurate improvements to our study of dynamical friction in isolation. Even so, we attempt to ameliorate errors from these choices by selecting galaxies by stellar mass throughout our analysis, which is significantly more stable against stripping than subhalo mass (see e.g. Peñarrubia et al., 2008; Smith et al., 2013, 2016). More importantly, any decrease in the subhalo mass will reduce the impact of dynamical friction, meaning that our results can be seen as an upper limit on how significantly DF impacts the splashback radius.
Our comparisons to other works (e.g. observations and cosmological simulations) of the splashback radius should be read with these limitations in mind. That our phase-space resembles that of IllustrisTNG (see Section 3.1) is a reassuring sign that we can compare our results to these other studies. Splashback radius studies typically stack clusters over redshift bins corresponding to spacings of $\gtrsim 2-3$ Gyr (see e.g. those of More et al., 2016; Chang et al., 2018; Murata et al., 2020), and furthermore, find that the splashback radius scales with $R_{\mathrm{200,mean}}$ across those redshift ranges (Chang et al., 2018). It is thus reasonable to compare our results with these studies despite the forward evolution of our simulation.
4.2 Residual trends
We find a significant residual trend towards smaller splashback radius as galaxy mass increases, even without dynamical friction (see Figure 5). The source of this trend is unclear, but it impacts primarily the lowest-mass clusters. Because our fit quality for the bootstrapped profiles in this mass bin is in fact better than that of higher mass clusters (see the discussion in Appendix A), we do not believe it results from a fitting issue. It is possible that the high-mass galaxies trace a different dynamical path than other galaxies, which appears more prominently among the lower-mass clusters.
Our galaxy samples are also greatly affected by the output of the TNG clusters at $z=0$, as they form the initial conditions of our simulations. O’Neil et al. (2022) worked with the same sample of haloes and subhaloes, and found similar, but larger, trends at $z=0$ (0 Gyr). That the magnitude of the residual trend we measure has decreased relative to O’Neil et al. (2022) suggests that the trend may arise from a non-DF process we do not model (e.g. baryonic physics).
We have examined the “contamination” of our sample at 5 Gyr (i.e. what fraction of galaxies within $2R_{\mathrm{200,mean}}$ at 5 Gyr were inside that radius at the start of our simulation), to see how these galaxies, which primarily trace the splashback radius, might be more or less affected by the starting conditions. However, this contamination fraction is relatively constant across cluster masses. Consequently, we cannot conclude that the $R_{\mathrm{sp}}$ measurements of lower mass clusters are differently impacted by the simulation starting conditions than high mass clusters. Instead, we conclude that massive galaxies trace a different path in the phase-space, particularly in low-mass clusters. We do not explore this difference further, though it is deserving of more study in the context of the splashback radius (see e.g. Rhee et al., 2017; Sampaio et al., 2024, for more discussion on galaxy mass and phase-space).
4.3 Dynamical friction implementation
The use of these idealized simulations, and the analytical implementation of dynamical friction, allows us to uniquely isolate DF. However, the analytical implementation of DF includes approximations not present in an N-body approach. Appendix B introduces a limited parameter study of the Coulomb logarithm $\Lambda$, an uncertain term in Chandrasekhar’s analytical DF formalism. We find that the uncertainties in the Coulomb logarithm do not significantly affect the location of the splashback radius (measured by subhalo positions). High-mass clusters are even less responsive to adjustments in $\Lambda$.
We also note that Chandrasekhar’s dynamical friction formula is thought to overestimate the role of DF compared to N-body simulations (see the merging timescales found by Boylan-Kolchin et al., 2008; Jiang et al., 2008), again suggesting that our results are an upper limit on the strength of dynamical friction.
4.4 Subhalo mass fraction requirement
In the literature, a variety of approaches, typically in simulation, have been used to find the minimum mass fraction ($M_{\rm subhalo}/M_{\rm halo}$) that is required for dynamical friction to have a significant impact on the subhalo orbit, which we expect will correspond to a change in the splashback radius. Though the subhalo mass is the dynamically relevant quantity, we express our results in terms of stellar mass because it is far more relevant for observations.
@@ 4.4 Subhalo mass fraction requirement
Work by van den Bosch et al. (1999) performed N-body simulations of an isothermal halo with Milky Way-like mass $\sim 10^{12}~\mathrm{M_{\odot}}$, and studied eccentricities of orbiting objects. They found DF timescales to be Gyr-scale, and that DF is negligible for satellites with mass under $10^{9}~\mathrm{M_{\odot}}$, producing a mass fraction requirement of $M_{\mathrm{subhalo}}/M_{\mathrm{halo}}~>~0.1\%$. They argue that this 0.1% figure scales for galaxy clusters, while assuming that cluster timescales are equivalent to the Milky Way simulation timescale. Therefore they expect dynamical friction to matter only for subhaloes of total mass $10^{12}~\mathrm{M_{\odot}}$, within a hypothetical cluster of mass $10^{15}~\mathrm{M_{\odot}}$. A subhalo mass of $10^{12}~\mathrm{M_{\odot}}$ corresponds to a stellar mass of $\sim 2\times 10^{10}~\mathrm{M_{\odot}}$ (see Table 2). Using a cut of $M_{*}>10^{10}~\mathrm{M_{\odot}}$ for our most massive cluster bin, we do not see a significant reduction in $R_{\mathrm{sp}}$, suggesting that a mass fraction above 0.1% may be required.
More recent work by van den Bosch et al. (2016), however, argues that a mass ratio of 10% is necessary for dynamical friction to notably alter an orbit, among haloes of mass $\gtrsim 10^{13}~\mathrm{M_{\odot}}$. They examine three different cosmological simulations (from the Bolshoi and Chinchilla suites) with varying cosmological parameters and mass resolutions. They find this mass fraction requirement across all three simulations. The 10% figure is consistent for our findings in the high cluster-mass regime. For example, among our most massive clusters, ($\sim 10^{15}~\mathrm{M_{\odot}}$) we would expect DF to be impactful only for stellar mass cuts exceeding $7\times 10^{11}~\mathrm{M_{\odot}}$. We thus would not expect to find significant differences for any of the stellar mass cuts shown in Figure 5, and we do not. On the other end, for a cluster of mass $10^{13}~\mathrm{M_{\odot}}$ in our sample, we would require a minmum stellar mass of $\sim 2\times 10^{10}~\mathrm{M_{\odot}}$ for DF to have an impact. We do see a notable ($\gtrsim 1\sigma$) reduction in the splashback radius for a cut of $M_{*}>10^{10}~\mathrm{M_{\odot}}$ after 5 Gyr, suggesting that the requirement of van den Bosch et al. (2016) may be too stringent. However, direct comparison is challenging because they did not divide their sample by halo mass. They, additionally, account for processes such as tidal stripping that contribute to subhalo mass loss, which are excluded from our study. Consistent with our comparison, it is expected that a higher starting mass fraction would be required if mass loss processes are included, because mass loss should reduce the strength of dynamical friction.
Yet other studies have sought a middle ground, finding that a mass fraction of 1% is a suitable requirement for dynamical friction to have an impact on subhaloes (Adhikari et al., 2016; More et al., 2016). Adhikari et al. (2016) examined clusters of virial mass $\geq 10^{14}$, and subhaloes with radial plunging orbits. They found the splashback radius was reduced by up to $\sim 20\%$ for low-mass clusters ($3-5\times 10^{13}~\mathrm{M_{\odot}}$) between subhaloes with masses less than- versus exceeding $4\times 10^{11}~\mathrm{M_{\odot}}$, which corresponds to a stellar mass cut of $\sim 3\times 10^{9}~\mathrm{M_{\odot}}$. In the matching cluster mass bin ($10^{13.5}\leq M_{\mathrm{200,mean}}/\mathrm{M_{\odot}}<10^{14}$), we find only a 7% reduction in the splashback radius, even using a stellar mass cut of $10^{10}~\mathrm{M_{\odot}}$. Plunging radial orbits (as used in Adhikari et al., 2016) should introduce the most extreme case of DF, so it is unsurprising that their subhaloes are more affected than those of this simulation. They also find that DF has a reduced effect in lower-concentration clusters, bringing their results closer to ours. Future work could examine the role of cluster concentration with the IllustrisTNG sample, using our more physically motivated subhalo orbits.
The wide variation in literature mass fractions reinforces the need for numerical simulations to ascertain the complex dependence of DF on subhalo mass, size, position, velocity, as well as halo density. We generally find that requiring $M_{\mathrm{subhalo}}/M_{\mathrm{halo}}~>~1\%$ captures our results, though again, we argue that $M_{\mathrm{*}}/M_{\mathrm{halo}}$ can be a more useful metric. For clusters of mass greater than $10^{14}\mathrm{M_{\odot}}$, a reduced splashback radius should be traced only by galaxies with stellar masses exceeding $10^{10}$ M, consistent with Figure 5.
4.5 Splashback radius measurements